# What Media Frames Reveal About Stance: A Dataset and Study about Memes in Climate Change Discourse

### v2

@@ Abstract
Media framing refers to the emphasis on specific aspects of perceived reality to shape how an issue is defined and understood. Its primary purpose is to shape public perceptions often in alignment with the authors’ opinions and stances. However, the interaction between stance and media frame remains largely unexplored. In this work, we apply an interdisciplinary approach to conceptualize and computationally explore this interaction with internet memes on climate change. We curate ClimateMemes, the first dataset of climate-change memes annotated with both stance and media frames, inspired by research in communication science. ClimateMemes includes 1,184 memes sourced from 47 subreddits, enabling analysis of frame prominence over time and communities, and sheds light on the framing preferences of different stance holders. We propose two meme understanding tasks: stance detection and media frame detection. We evaluate LLaVA-NeXT and Molmo in various setups, and report the corresponding results on their LLM backbone. Human captions consistently enhance performance. Synthetic captions and human-corrected OCR also help occasionally. Our findings highlight that VLMs perform well on stance, but struggle on frames, where LLMs outperform VLMs. Finally, we analyze VLMs’ limitations in handling nuanced frames and stance expressions on climate change internet memes.
@@ A Dataset and Study about Memes in Climate Change Discourse
[…]
@@ 1 Introduction
Memes are a powerful communication format in online discourse that reflect communities’ cultural and social dynamics  Davis et al. (2016); Zhang and Pinto (2021). Memes are multimodal digital items that combine images and texts to convey complex viewpoints in a compact and engaging format Sharma et al. (2020); Liu et al. (2022). Memes express the communicator’s position towards a particular topic, i.e., stances as Mohammad et al. (2016) defines, through the specific combination of modalities such as text and images.
While stance reflects the creator’s opinion toward a target, the specific narrative used to convey a certain stance is shaped by media frames. Media framing refers to selecting specific aspects of a perceived reality in communication to portray how an issue is defined, how its causes are interpreted, how its moral implications are evaluated, and what potential solutions are considered Entman (1993); Gidin (1980). Depending on their stance, creators may gravitate toward different framing strategies Snow and Benford (1992). However, the interaction between stance and media frames remains under-studied, particularly in their representation through humorous social media content, such as memes, and its impact on debates of global significance, such as the climate change.
Climate change (CC) memes are a vital component of social media, such as Twitter/X Ross and Rivers (2019). For example, Figure 1(a) conveys a convinced stance towards CC by using and impact frames (further detailed in §3.2) to affirm the evidence of global warming and its disheartening consequences. Conversely, Figure 1(b) conveys a skeptical stance using the hoax frame, claiming that CC is not a major issue or even not real and suggests that politics may distort the CC issue.
In this paper, we analyze stances and media frames in CC memes111 We will release the data and code with the camera-ready submission. by examining the following three research questions (RQs):
@@
  • RQ2: Can state-of-the-art VLMs and LLMs accurately detect stances conveyed by memes and the corresponding media frames? We extend stance detection from text
    and propose a new task of multi-label media frame detection on CC memes. We evaluate two open-source VLMs and their backbone LLMs (§5.1) and investigate the effects of few-shot experiments and input modalities on these two tasks (§5.2). We found that while synthetic meme captions cannot yet fully replace human-annotated ones, they still improve the VLMs’ performance on both tasks. Yet, LLMs outperform VLMs on frame detection.
  • RQ3: Can taxonomies from communication science provide more insights on stance and media frame detection results? We recruit communication science specialists to annotate humor type, person, and responsibility features on 235 test CC memes. Our analyses reveal that the performances of VLMs and LLMs degrade markedly on memes that are jokes, about political figures, and individual (micro-level) responsibilities (§6).
@@ 2.1 Memes
Nguyen and Ng (2024) categorize meme understanding into three main types: classification, interpretation, and explanation. Classification aims to assign labels to memes, such as identifying harmful content  Kiela et al. (2020); Pramanick et al. (2021); Cao et al. (2022); Hee et al. (2023), sentiment  Sharma et al. (2020); Chauhan et al. (2020), or figurative language Liu et al. (2022); Xu et al. (2022). Interpretation tasks focus on understanding and generating insights from memes, such as generating captions or analyzing the metaphor between the image and text componants Hwang and Shwartz (2023); Chen et al. (2024). Explanation tasks go a step further by generating textual justifications for the labels assigned to memes (Hee et al., 2023). In this study, we curate ClimateMemes and introduce two meme understanding tasks: stance detection and media frame detection. We also collect human-corrected OCR and meme captioning as a basis for future tasks.
@@ 2.2 Media Frames
In the context of climate change, framing has been studied to understand its role in public discourse and media representation. Stede et al. (2023) utilize generic frames, which are more abstract and commonly observed across political discussions, to analyze climate change in Nature and Science editorials. Chen et al. (2022) study how frames evolve within public events, emphasizing their divergence and convergence in shaping climate change narratives, while Frermann et al. (2023) analyze how news articles across the political spectrum frame climate change. To the best of our knowledge, this paper presents the first dataset of multimodal memes annotated with media frames and analyzes how frames interact with stances.
@@ 3 ClimateMemes Dataset
This section describes ClimateMemes, a dataset of 1,184 CC memes from 47 subreddits annotated with media frames and stances. Figure 2 illustrates our data processing pipeline. We discuss meme collection and climate filtering (§3.1), and present guidelines for stance and frame annotations (§3.2).
@@ 3.1 Source Memes and Climate Filter
Data Source.
To collect CC memes, we search subreddits with “meme” in their names and filter the topic of posts with the keyword “climate”. To obtain diverse climate change perspectives, our collection includes subreddits like r/ClimateMemes (primarily hosting climate activists) and r/ConservativeMemes (reflecting a more skeptical community on CC).
@@ 4.1 ClimateMemes Statistics
Despite continuous efforts to upsample skeptical memes and subreddits, ClimateMemes exhibits an unbalance where 78.0% memes are convinced and 17.2% are skeptical, most frequently employing adequacy and hoax frames, respectively (see Appendix C for a detailed frame distribution).
@@ 4.2 Frame Preference
Figure 4 shows the probability of particular frames being involved in memes with convinced and skeptical stances. Among skeptical memes, 77.94% involve hoax, followed by propriety at 45.59%. Other frames appear in less than 15% memes. In contrast, frames in convinced memes are more diverse, with adequacy, impact, and cause being the most common, appearing in 42.1%, 40.20%, and 37.05% memes. Other frames, except for prospect, appear in 20%-30% memes.
@@ 4.4 Frame as a signal
Given that specific frames such as hoax are prevalent in skeptical memes, we examine whether frames serve as a good signal for stance detection. Figure 6 analyzes the likelihood of a meme being convinced or skeptical when a specific frame is used. We observe that when cause, impact, adequacy, and prospect appear in a meme, there is $>$80% probability that the meme holds a convinced stance. and allocation also appear more frequently in convinced memes. Conversely, hoax implies a 76.18% probability that the meme is skeptical, followed by propriety (59.87%).
@@ 5.1 Experimental Setups
Models.
We evaluate two open-source VLMs on multimodal memes: LLaVA-v1.6-Mistral-7B (LLaVA, Liu et al. 2024), Molmo-7B-D (Molmo,  Deitke et al. 2024), both of a visual encoder→cross-modal connector→LLM setup.
To compare, we experiment with text-only inputs on their LLM backbones: Mistral-7B (Mistral, Jiang et al. 2023) and Qwen2-7B (Qwen, Yang et al. 2024).
@@
Evaluation Scenarios.
In addition to zero-shot, we evaluate all models on $n$-shot experiments $n$ ranging from 1 to 4. Following Huang et al. (2024), we leverage relative sample augmentation to select top $n$ similar memes from train for each test meme based on the image and its human-corrected OCR. We also explore various input scenarios following Hwang and Shwartz (2023) to examine whether OCR, human, and synthetic meme caption can improve stance and media frame detection. We rotate stance and frame orders in prompts and report average over permutations (Zheng et al., 2023; Wang et al., 2024). For the backbone LLM baselines, we run experiments on text-only inputs.
@@
Acc. F1 Acc. F1
baseline meme 80.85 29.80 73.83 43.98
LLaVA meme 77.31 39.08 51.87 45.63
meme+ocr 77.31 44.06 46.36 40.72
meme+syn 73.95 40.01 52.45 45.78
meme+syn+ocr 76.89 41.10 52.57 45.87
meme+hum 86.55 56.68 49.96 44.18
meme+hum+ocr 83.19 53.57 50.53 44.46
Molmo meme 47.06 28.16 60.37 52.60
meme+ocr 57.56 34.70 56.98 49.68
meme+syn 61.76 39.25 58.37 51.02
meme+syn+ocr 65.97 38.32 54.23 47.97
meme+hum 72.27 49.53 62.74 54.24
meme+hum+ocr 70.17 46.52 60.40 52.46
Mistral ocr 51.90 37.09 61.71 54.79
syn 58.23 36.06 59.03 53.01
syn+ocr 59.66 42.71 61.78 55.20
hum 79.32 60.54 64.61 58.31
hum+ocr 67.65 48.96 65.09 58.78
Qwen ocr 49.16 34.06 64.02 55.45
syn 68.91 44.66 60.33 53.98
syn+ocr 61.34 39.08 60.88 54.24
hum 73.11 53.28 65.86 58.23
hum+ocr 70.17 51.66 64.98 57.51
@@ 5.2 Which inputs help stance and frame detection in memes?
Zero-shot vs. Few-shot.
@@
VLMs vs. LLM backbones.
To what extent can visual inputs benefit VLM performances on meme understanding? While LLaVA has an edge over Mistral across various inputs on stance detection, both VLMs achieve lower scores on frame detection compared to LLMs. We hypothesize that VLMs are not pre-trained on meme datasets for frame detection. Yet, there already exists textual dataset related to framing Stede and Patz (2021); Frermann et al. (2023). It should also be noted that LLMs’ winning performances benefit from costly human annotations (OCR corrections [^4] We observed low-quality OCR; the average Levenshtein edit-distance before and after human correction is 60.75. and captions) or synthetic captions generated by VLMs. We also find that certain visual elements can mislead VLMs. Figure 7 shows two examples where LLaVA over-associates fire with . Though wildfires are often associated with CC existence, flames may also appear as metaphor or background.
@@
OCR.
@@
Human vs. Synthetic Caption.
@@
Frame #M LLaVA Molmo Mistral Qwen
Acc. F1 Acc. F1 Acc. F1 Acc. F1
44 30.69 26.84 68.99 59.43 46.60 43.88 72.13 60.32
hoax 81 51.60 46.90 71.54 59.43 71.33 60.54 71.01 61.46
caus 75 49.47 45.90 64.63 57.06 70.37 60.64 65.85 58.22
impa 77 48.03 45.37 60.59 54.64 71.33 61.29 61.54 56.03
allo 49 52.55 48.05 60.00 55.22 69.95 60.17 64.36 56.99
prop 80 53.62 49.54 56.86 53.18 64.10 56.65 64.10 57.57
adeq 81 56.86 50.50 59.68 54.84 64.26 56.98 65.21 58.28
pros 13 56.86 50.69 59.63 54.74 58.99 53.58 62.71 56.99
average 49.96 45.47 62.74 56.07 64.61 56.72 65.86 58.23
@@ 5.3 Which frames are harder?
Table 3 reports per-frame performances of VLMs and LLMs. Consistent with overall performance, Molmo outperforms LLaVA in predicting all 8 frames. hoax achieves the highest score and propriety scores the lowest. For LLMs, Qwen outperforms Mistral with overall performance, but is not consistently the best model for all frames.
Model Base Input(+frame) $\Delta$Acc. $\Delta$F1
LLaVA meme -5.04 +0.45
meme+ocr -2.52 +5.26
meme+hum 0.00 +1.06
meme+hum+ocr -15.54 -7.03
meme+syn +2.52 +2.27
meme+syn+ocr -16.39 -4.74
Molmo meme +1.26 +2.81
meme+ocr +7.15 +3.33
meme+hum +5.46 +4.15
meme+hum+ocr +7.14 -0.30
meme+syn +6.31 +2.73
meme+syn+ocr +7.56 +3.11
@@ 5.4 Can frame labels help stance detection?
Table 4 investigates whether adding gold frame labels helps stance detection on 4-shot VLMs. Notably, for LLaVA (better at stance detection), incorporating frame information tends to reduce its performance. Instead, for Molmo (better at frame detection), adding frame information generally boosts its performance. Both models further improve their performance in the image and human caption setup. This suggests that stance and frame detection could benefit from multi-task training, improving performance through shared knowledge.

### v3

@@ Abstract
Media framing refers to the emphasis on specific aspects of perceived reality to shape how an issue is defined and understood. Its primary purpose is to shape public perceptions, often in alignment with the authors’ opinions and stances. However, the interaction between stance and media frame remains largely unexplored. In this work, we apply an interdisciplinary approach to conceptualize and computationally explore this interaction with internet memes on climate change. We curate ClimateMemes, the first dataset of climate-change memes annotated with both stance and media frames, inspired by research in communication science. ClimateMemes includes 1,184 memes sourced from 47 subreddits, enabling analysis of frame prominence over time and communities, and sheds light on the framing preferences of different stance holders. We propose two meme understanding tasks: stance detection and media frame detection. In various input setups, we evaluate two vision language models (VLMs), LLaVA-NeXT and Molmo, and report the corresponding results on their backbone large language models (LLMs). Human captions consistently enhance performance. Synthetic captions and human-corrected OCR also help occasionally. Our findings highlight that VLMs perform well on stance, but struggle on frames, where LLMs outperform VLMs. Finally, we draw on concepts from Computational Communication Science to analyze VLMs’ limitations, showing that memes employing specific humor types, personalization, and responsibility cues pose challenges for VLMs in handling nuanced frames and stance expressions on climate change.  [^1] We make our dataset publicly available at https://github.com/mainlp/ClimateMemes.
@@ A Dataset and Study about Memes in Climate Change Discourse
Shijia ZhouU*   Siyao PengU*   Simon M. Luebke Jörg Haßler   Mario Haim   Saif M. Mohammad2   Barbara PlankU U MaiNLP & MCML, LMU Munich, Germany Department of Media and Communication, LMU Munich, Germany 2 National Research Council Canada, Ottawa, Canada {zhou.shijia, siyao.peng, mario.haim, b.plank}@lmu.deAnonymous ACL submission
@@ 1 Introduction
Internet memes are a powerful communication format in online discourse that reflect communities’ cultural and social dynamics  Davis et al. (2016); Zhang and Pinto (2021). Multimodal digital items combine images and texts to convey complex viewpoints in a compact and engaging format Sharma et al. (2020); Liu et al. (2022). Through these multimodal expressions, communicators convey their positions towards particular topics, i.e., stances as defined by Mohammad et al. (2016).
While stance reflects the creator’s opinion toward a target, the specific narrative used to convey a certain stance is shaped by media frames. Media framing refers to selecting specific aspects of a perceived reality in communication to portray how an issue is defined, how its causes are interpreted, how its moral implications are evaluated, and what potential solutions are considered Entman (1993); Gidin (1980). Depending on their stance, creators may gravitate toward different framing strategies Snow and Benford (1992). However, the interaction between stance and media frames remains under-studied, particularly in their representation through humorous social media content such as memes. This is especially relevant for debates of global significance, such as climate change.
Memes about climate change (CC) are widespread on social media, including Twitter/X Ross and Rivers (2019). For example, Figure 1(a) conveys a convinced stance towards CC by using and impact frames (further detailed in §3.2) to affirm the evidence of global warming and its disheartening consequences. Conversely, Figure 1(b) conveys a skeptical stance using the hoax frame, claiming that CC is not a major issue or even not real, and suggests that politics may distort the CC issue.
In this paper, we analyze stances and media frames in CC memes by examining the following three research questions (RQs):
@@
  • RQ2: Can state-of-the-art VLMs and LLMs accurately detect stances conveyed by memes and the corresponding media frames? We extend stance detection from text and propose a new task of multi-label media frame detection on CC memes. We evaluate two open-source VLMs and their backbone LLMs (§5.1) and investigate the effects of few-shot experiments and input modalities on these two tasks (§5.2). We found that while synthetic meme captions cannot yet fully replace human-annotated ones, they still improve the VLMs’ performance on both tasks. Yet, LLMs outperform VLMs on frame detection.
  • RQ3: Can taxonomies from communication science provide more insights into stance and media frame detection results? We recruit communication science specialists to annotate humor type, person, and responsibility features on 235 test CC memes. Our analyses reveal that the performances of VLMs and LLMs degrade markedly on memes that are jokes, about political figures, and about individual (micro-level) responsibilities (§6).
@@ 2.1 Memes
Nguyen and Ng (2024) categorize meme understanding into three main types: classification, interpretation, and explanation. Classification aims to assign labels to memes, such as identifying harmful content  Kiela et al. (2020); Pramanick et al. (2021); Cao et al. (2022); Hee et al. (2023); Chen et al. (2025); Liu et al. (2025); Chakravarthi et al. (2025), sentiment  Sharma et al. (2020); Chauhan et al. (2020); Nguyen et al. (2025), or figurative language Liu et al. (2022); Xu et al. (2022). Interpretation tasks focus on understanding and generating insights from memes, such as generating captions or analyzing the metaphor between the image and text componants Hwang and Shwartz (2023); Chen et al. (2024). Explanation tasks go a step further by generating textual justifications for the labels assigned to memes (Hee et al., 2023). In this study, we curate ClimateMemes and introduce two meme understanding tasks: stance detection and media frame detection. We also collect human-corrected OCR and human-written meme captioning as a basis for future tasks.
@@ 2.2 Media Frames
In the context of climate change, framing has been studied to understand its role in public discourse and media representation (Otmakhova and Frermann, 2025). Stede et al. (2023) utilize generic frames, which are more abstract and commonly observed across political discussions, to analyze climate change in Nature and Science editorials. Chen et al. (2022) study how frames evolve within public events, emphasizing their divergence and convergence in shaping climate change narratives. Frermann et al. (2023) analyze how news articles across the political spectrum frame climate change. To the best of our knowledge, this paper presents the first dataset of multimodal memes annotated with media frames and analyzes how frames interact with stances.
@@ 3 ClimateMemes Dataset
This section describes ClimateMemes, a dataset of 1,184 CC memes from 47 subreddits annotated with media frames and stances. Figure 2 illustrates our data processing pipeline. We discuss meme collection and climate filtering (§3.1), and present guidelines for stance and frame annotations (§3.2). We also provide manual OCR correction and meme caption annotation for future uptake (§3.3).
@@ 3.1 Source Memes and Climate Filter
Data Source
To collect CC memes, we search subreddits with “meme” in their names and filter the topic of posts with the keyword “climate.” To obtain diverse climate change perspectives, our collection includes subreddits like r/ClimateMemes (primarily hosting climate activists) and r/ConservativeMemes (reflecting a more skeptical community on CC).
@@ 4.1 ClimateMemes Statistics
Despite continuous efforts to upsample skeptical memes and subreddits, ClimateMemes exhibits an imbalance where 78.0% memes are convinced and 17.2% are skeptical, most frequently employing adequacy and hoax frames, respectively (see Appendix C for a detailed frame distribution).
@@ 4.2 Frame Preference
Figure 4 shows the probability of particular frames being involved in memes with convinced and skeptical stances. Among skeptical memes, 77.94% involve hoax, followed by propriety at 45.59%. Other frames appear in less than 15% of memes. In contrast, frames in convinced memes are more diverse, with adequacy, impact, and cause being the most common, appearing in 42.1%, 40.20%, and 37.05% of memes. Other frames, except for prospect, appear in 20%-30% of memes.
@@ 4.4 Frame as a signal
Given that specific frames such as hoax are prevalent in skeptical memes, we examine whether frames serve as a good signal for stance detection. Figure 6 analyzes the likelihood of a meme being convinced or skeptical when a specific frame is used. We observe that when cause, impact, adequacy, and prospect appear in a meme, there is >>80% probability that the meme holds a convinced stance. and allocation also appear more frequently in convinced memes. Conversely, hoax implies a 76.18% probability that the meme is skeptical, followed by propriety (59.87%).
@@ 5.1 Experimental Setups
Models
We evaluate two open-source VLMs on multimodal memes: LLaVA-v1.6-Mistral-7B (LLaVA, Liu et al. 2024) and Molmo-7B-D (Molmo,  Deitke et al. 2024), both of a visual encoder→cross-modal connector→LLM setup. To compare, we experiment with text-only inputs on their LLM backbones: Mistral-7B (Mistral, Jiang et al. 2023) and Qwen2-7B (Qwen, Yang et al. 2024).
@@
Evaluation Scenarios
In addition to zero-shot, we evaluate all models on $n$-shot experiments $n$ ranging from 1 to 4. Following Huang et al. (2024), we leverage relative sample augmentation to select top $n$ similar memes from train for each test meme based on the image and its human-corrected OCR. We also explore various input scenarios following Hwang and Shwartz (2023) to examine whether manually-corrected OCR and human meme caption (detailed in §3.3) can improve stance and media frame detection, as well as synthetic caption generated by VLMs. We rotate stance and frame orders in prompts and report the average over permutations (Zheng et al., 2023; Wang et al., 2024). For the backbone LLM baselines, we run experiments on text-only inputs.
@@
F1 Acc. F1 Acc.
baseline meme 29.80 80.85 43.98 73.83
LLaVA meme 39.08 77.31 45.63 51.87
meme+OCR 44.06 77.31 40.72 46.36
meme+syn 40.01 73.95 45.78 52.45
meme+syn+OCR 41.10 76.89 45.87 52.57
meme+hum 56.68 86.55 44.18 49.96
meme+hum+OCR 53.57 83.19 44.46 50.53
Molmo meme 28.16 47.06 52.60 60.37
meme+OCR 34.70 57.56 49.68 56.98
meme+syn 39.25 61.76 51.02 58.37
meme+syn+OCR 38.32 65.97 47.97 54.23
meme+hum 49.53 72.27 54.24 62.74
meme+hum+OCR 46.52 70.17 52.46 60.40
Mistral OCR 37.09 51.90 54.79 61.71
syn 36.06 58.23 53.01 59.03
syn+OCR 42.71 59.66 55.20 61.78
hum 60.54 79.32 58.31 64.61
hum+OCR 48.96 67.65 58.78 65.09
Qwen OCR 34.06 49.16 55.45 64.02
syn 44.66 68.91 53.98 60.33
syn+OCR 39.08 61.34 54.24 60.88
hum 53.28 73.11 58.23 65.86
hum+OCR 51.66 70.17 57.51 64.98
@@ 5.2 Which inputs help stance and frame detection in memes?
Zero-shot vs. Few-shot
@@
VLMs vs. LLM backbones
To what extent can visual inputs benefit VLM performances on meme understanding? While LLaVA has an edge over Mistral across various inputs on stance detection (with the exception of meme+syn+OCR and meme+hum for VLMs), both VLMs achieve lower scores on frame detection compared to LLMs. We hypothesize that VLMs are not pre-trained on meme datasets for frame detection. Yet, there already exists textual dataset related to framing Stede and Patz (2021); Frermann et al. (2023). It should also be noted that LLMs’ winning performances benefit from costly human annotations (OCR corrections [^4] We observed low-quality OCR; the average Levenshtein edit-distance before and after human correction is 60.75. and captions) or synthetic captions generated by VLMs.
Through qualitative analyses, we find that certain visual elements can mislead VLMs. Figure 7 shows two examples where LLaVA over-associates flames in image with . The template of Figure 7(a) comes from the movie Thor: Ragnarok, where the flames on the mountain are used to dramatize a war scene. Figure 7(b) satirizes the narrative that attributes all environmental issues – such as forest fires caused by arson – to climate change. Although flames in our dataset are often associated with the existence of climate change, in these cases they do not depict actual climate change disasters but instead serve as background elements or as examples used to cast doubt on climate change.
@@
Frame #M LLaVA Molmo Mistral Qwen
F1 Acc. F1 Acc. F1 Acc. F1 Acc.
44 26.84 30.69 59.43 68.99 43.88 46.60 60.32 72.13
hoax 81 46.90 51.60 59.43 71.54 60.54 71.33 61.46 71.01
caus 75 45.90 49.47 57.06 64.63 60.64 70.37 58.22 65.85
impa 77 45.37 48.03 54.64 60.59 61.29 71.33 56.03 61.54
allo 49 48.05 52.55 55.22 60.00 60.17 69.95 56.99 64.36
prop 80 49.54 53.62 53.18 56.86 56.65 64.10 57.57 64.10
adeq 81 50.50 56.86 54.84 59.68 56.98 64.26 58.28 65.21
pros 13 50.69 56.86 54.74 59.63 53.58 58.99 56.99 62.71
[…]
OCR
@@
Human vs. Synthetic Caption
@@ 5.3 Which frames are harder?
Table 3 reports per-frame performances of VLMs and LLMs. Consistent with overall performance, Molmo outperforms LLaVA in predicting all 8 frames. Molmo scores the highest on hoax and the lowest on propriety. For LLMs, Qwen outperforms Mistral with overall performance, but is not the best in every frame.
[…]
meme+hum +1.06 0.00
[…]
meme+syn +2.27 +2.52
[…]
Molmo meme +2.81 +1.26
meme+OCR +3.33 +7.15
meme+hum +4.15 +5.46
[…]
meme+syn +2.73 +6.31
meme+syn+OCR +3.11 +7.56
@@ 5.4 Can frame labels help stance detection?
Table 4 investigates whether adding gold frame labels helps stance detection on 4-shot VLMs. Notably, for LLaVA (better at stance detection), incorporating frame information leads to F1 drops of 7.03 with meme+hum+OCR and 4.74 with meme+syn+OCR, while the other four setups show improvements. Instead, for Molmo (better at frame detection), adding frame information generally boosts its performance. Both models further improve their performance in the image and human caption setup. This suggests that stance and frame detection could benefit from multi-task training, improving performance through shared knowledge.